# DLCR: A Generative Data Expansion Framework via Diffusion for Clothes-Changing Person Re-ID

### v1

@@ Abstract
With the recent exhibited strength of generative diffusion models, an open research question is if images generated by these models can be used to learn better visual representations. While this generative data expansion may suffice for easier visual tasks, we explore its efficacy on a more difficult discriminative task: clothes-changing person re-identification (CC-ReID). CC-ReID aims to match people appearing in non-overlapping cameras, even when they change their clothes across cameras. Not only are current CC-ReID models constrained by the limited diversity of clothing in current CC-ReID datasets, but generating additional data that retains important personal features for accurate identification is a current challenge. To address this issue we propose DLCR, a novel data expansion framework that leverages pre-trained diffusion and large language models (LLMs) to accurately generate diverse images of individuals in varied attire. We generate additional data for five benchmark CC-ReID datasets (PRCC, CCVID, LaST, VC-Clothes, and LTCC) and increase their clothing diversity by $10$x, totaling over $2.1$M images generated. DLCR employs diffusion-based text-guided inpainting, conditioned on clothing prompts constructed using LLMs, to generate synthetic data that only modifies a subject’s clothes while preserving their personally identifiable features. With this massive increase in data, we introduce two novel strategies - progressive learning and test-time prediction refinement - that respectively reduce training time and further boosts CC-ReID performance. We validate our method through extensive ablations and experiments and show massive improvements when training previous CC-ReID methods with our generated data. On the PRCC dataset, we obtain a large top-1 accuracy improvement of $11.3\%$ by training CAL, a previous state of the art (SOTA) method, with DLCR-generated data. We publicly release our code and generated data for each dataset here: https://github.com/CroitoruAlin/dlcr.
@@ 1 Introduction
A majority of computer vision research can be categorized as discriminative methods, wherein improved or novel model architectures are proposed and shown to learn better visual representations than their predecessors. These methods operate under the assumption that their training datasets are sufficiently large and diverse, with the limiting factor being the model’s ability to learn the desired objective. However, when there is a lack in diversity in the commonly used datasets for some task, the limiting factor in this case becomes the training data itself. Collecting additional in-domain data can be both difficult and expensive, and only recently have generative models become viable enough to generate additional, high-fidelity training data (which we refer to as generative data expansion). Generative data expansion has been previously explored for supervised or self-supervised learning on easier discriminative tasks, such as ImageNet classification [3, 50]. While there are other unexplored, more difficult discriminative visual tasks that may benefit even more so from generative data expansion, the difficulty lies in generating accurate, in-domain data. In this work, we tackle this problem in the context of CC-ReID, which is difficult in nature and, due to its relative infancy, a majority of datasets suffer from only containing a small number ($2-5$) of clothing outfits per subject [52, 58, 39, 15, 45].
@@
In Person Re-identification [60, 61], the primary objective is to match an individual across images or videos captured by cameras at different times and locations. This task becomes notably challenging due to disparities in the body poses, lighting, backgrounds, viewpoints, and potential occlusions present across cameras. Traditional Re-ID approaches [5, 8, 38], also known as short-term Re-ID, assume that individuals are observed for relatively brief durations and that their attire remains constant throughout videos. However, these assumptions often fall short of practical real-world scenarios. For instance, people’s clothes can change when captured across different camera feeds at different times and days. Consequently, a new branch of Re-ID, known as clothes-changing Re-ID [39, 15, 24, 17], has recently emerged to reliably identify people even when their clothing constantly change over extended periods of time. Recent CC-ReID works commonly make discriminative modifications to traditional Re-ID model architectures and loss formulations to disentangle structural (clothes-independent) and appearance information (clothes-dependent) in the learned feature space. However, these discriminative approaches are inherently constrained by the aforementioned low clothing variations present in current CC-ReID datasets.
@@
Figure 1: Best viewed with zoom. (Left) Comparing alternative data expansion approaches to our method. Simply changing the color of the clothes using our clothing masks preserves the subject’s ID information, but lacks diversity. Using standard diffusion without inpainting can sometimes introduce diversity, but the ID-related information in an image is destroyed. Our method strongly increases clothing diversity, even with specific and difficult prompts, and preserves the subject’s ID-related information for effective CC-ReID. (Right) DLCR significantly increases the size of training data for five benchmark CC-ReID datasets, generating over $2.1$M diverse, clothes-changed images.
@@
To address these issues, in this work we propose an effective data expansion framework for CC-ReID (named DLCR) that increases the number of clothing variations in CC-ReID datasets by intelligently harnessing generative AI and foundation models (Diffusion and large Language models for CC-ReID). DLCR consists of two stages: (1) Data Generation and (2) Re-ID Training and Prediction Refinement. In the first stage, our objective is to enrich a given CC-ReID training dataset by generating additional samples where clothing items are artificially changed through inpainting. With more diverse training images of a particular person in different clothes, a CC-ReID model will learn better clothes-invariant person features during training. To achieve this objective, we utilize diffusion models [20, 42, 6, 36], which are capable of generating realistic images and inpainting for image editing. However, diffusion models often struggle with preserving the intricate details of a person’s identity during image generation (Fig. 1). We overcome this by leveraging a human parsing method [26] to produce binary masks that mark only the clothing regions in an image. We then use this binary mask to retain ID-specific portions in an image during diffusion inpainting, such as the face, hair, and body shape, thereby only augmenting the subject’s clothes. Furthermore, existing CC-ReID datasets only use scalar values for each unique clothes ID and lack textual descriptions of specific clothing items, preventing the use of text-guided diffusion models for data generation. Thus, to overcome this issue, we generate explicit clothing descriptions by jointly leveraging a large visual-language model (LLaVA [32]) and a large language model (LLaMA [51]).
@@
In the second stage of DLCR, we propose two effective strategies in which our generated data can be sophisticatedly utilized during CC-ReID training and testing. Firstly, we train various CC-ReID models using our generated data from stage 1 and show how our enriched training data improves performance in clothes-changing settings. To accommodate both the large increase in data and the higher clothing variety of the generated samples, we employ a progressive learning strategy during training, where we gradually introduce the generated clothing variations of a person. This allows for training on $10$x more data without drastically increasing training time, as well as further assists the model in learning better clothes-invariant person features. During inference, we apply our prediction refinement strategy which is specifically tailored for a CC-ReID retrieval task. We create different variations of a given query by artificially changing the subject’s attire, then ensemble their retrieved similarity scores from the gallery to obtain a refined final prediction.
@@ 2 Related Works
In standard Re-ID, clothing is considered a static attribute, which leads standard Re-ID models to erroneously couple clothing items with specific subjects. CC-ReID is a more difficult paradigm that requires models to intrinsically learn person-specific features that are robust to covariate information, especially clothing items. To achieve this robustness, both discriminative [19, 16, 21, 58, 28, 11, 62, 31, 56, 18] and generative [35, 55, 13, 34, 63, 57] approaches for CC-ReID have been explored. Discriminative approaches focus on deliberate model architectures [30, 15] and loss formulations [19, 16, 21], while generative methods aim to synthesize data that increases the quality or diversity of an existing training set. The common goal of these methods is to enforce the learning of clothes-invariant person features, thereby improving CC-ReID performance on datasets in which subjects are captured in multiple different clothing outfits.
@@
Generative methods: With the aforementioned lack of such diverse datasets, the previously dominant approach has been to use Generative Adversarial Networks (GANs) [14] to synthesize additional training data for specific learning objectives, such as pose-invariance [35, 55, 13], clothes-invariance [34, 63, 57], and improving general performance [64, 29, 23]. However, these GAN-based works also suffer from constraints such as limited diversity/mode collapse [63, 33] and the limited capability of only transferring existing styles from one subject to another [9, 65, 66]. Most important to note, GANs are dataset-specific, meaning they require separate training on each dataset before data generation. Furthermore, one of the most successful GAN-based works [63] does not directly use its generated data during training, but rather transfers its dataset-specific generator encoder to the discriminator to produce better clothes-invariant person features, reducing generalizability. On the other hand, diffusion models [36] have recently overtaken GANs in a multitude of other vision domains [10, 6, 43, 44] as a superior class of generative models, yet have not been applied to the CC-ReID domain. Furthermore, the previously discussed limitations of GANs can be overcome by large-scale, pre-trained diffusion models which have shown impressive zero-shot generation abilities on a multitude of out-of-domain tasks [40, 42, 37]. DLCR is the first generative approach to harness the benefits of diffusion models for CC-ReID, and consequently, uses the same generator for every dataset. In a novel fashion, we leverage a text-conditioned diffusion inpainting model to synthesize completely new clothing outfits onto any subject in a dataset, separating the process between data generation and discriminative CC-ReID training.
@@ 3 Proposed Approach
DLCR is a generic data expansion method that can be applied to any CC-ReID dataset, wherein the given training data is enriched to improve data diversity for better learning of clothes-invariant features. Our method contains two stages: Data Generation (Sec. 2) and Re-ID Training and Prediction Refinement (Sec. 3.2). In stage 1, DLCR utilizes semantic human masks, a visual language model (LLaVA), a large language model (LLaMA), and text-conditioned diffusion inpainting to controllably change a subject’s clothes in an image. In stage 2, we leverage our generated data, in conjunction with our progressive learning strategy, to train CC-ReID models with $10$x more data with negligible impacts to training time. We also propose a test-time prediction refinement strategy specifically tailored for CC-ReID retrieval. Before explaining the method in detail, we provide a brief background on diffusion models.
@@ 3.1 Stage 1: Data Generation
In CC-ReID, the training dataset $D_{train}=\{(x_{i},s_{i},c_{i})\}_{i=1}^{M}$, is a collection of $M$ samples where an $i^{th}$ sample is a triplet consisting of an RGB image $x_{i}$, subject (person) ID $s_{i}$, and clothes ID $c_{i}$. Generating images of a particular subject, $s_{i}$, in different clothing outfits can enrich the training data by improving its diversity. Specifically, in this stage, we generate multiple variants of a training image, $x_{i}$, by artificially inpainting the clothes of subject $s_{i}$. We achieve this objective by using a text-conditioned diffusion inpainting model (Stable Diffusion Inpainting [42]), where text prompts describe the outfits to be generated in the clothing regions specified by a human-parsed semantic mask. These components (also shown in Fig. 2) are described below.
@@
Semantic Mask & Text Extraction: While generating different images of a subject, it is important to preserve the person-relevant information. However, diffusion models struggle to retain this information due to the varied and complex nature of human faces and body shapes. Thus, we apply a semantic human parsing method [26] to each image, $x_{i}$, to obtain an ID-preserving binary mask, $m_{i}$, where only the upper-clothes, lower-clothes, and footwear regions are marked for inpainting. This process is computationally negligible, with throughput as high as 60 masks per second. This allows the person ($s_{i}$) features, such as the face, hair, and general body structure, to be preserved in the generated samples, even after clothes inpainting; an especially critical aspect for enriching CC-ReID data. In general, the set of images belonging to a clothes ID, $c$, can be represented by $X^{c}=\{x^{c}_{i}\}_{i=1}^{n}$. After applying ID-preserving semantic masks to these images, we obtain a set of masked images $\bar{X}^{c}=\{x^{c}_{i}\circ m_{i}\}_{i=1}^{n}$. While CC-ReID datasets provide clothes IDs, they are usually simple scalar values that do not contain descriptive information as to what specific clothing items are present in an image. To extract these additional clothing descriptions for each clothes ID in $D_{train}$, we utilize LLaVA [32] and LLaMA [51] (see Fig. 2). A naive approach would be to simply generate descriptions of clothing items using an LLM, e.g. LLaMA, or to create random clothing descriptions. While this is likely to increase the diversity of the dataset, and consequently, the generalization capacity of the downstream CC-ReID models, it does not alleviate dataset-specific biases. We use LLaVA in order to obtain descriptions of clothing items that are present in the dataset, aiming to reduce the dataset-specific bias (see supplementary). This forces the CC-ReID models to focus on identity features, and ignore clothing features.
@@
Diffusion-Based Clothes Inpainting: A text-conditioned diffusion inpainting model, paired with our extracted clothing descriptions and semantic masks, can be used to controllably generate diverse, synthetic samples. To this end, we propose, for the first time, the use of a diffusion inpainting model for CC-ReID. We first construct a set $C=\{c_{i}\}_{i=1}^{L}$ containing all the unique clothes IDs in the training data $D_{train}$. For each training image $x_{i}$ with subject ID $s_{i}$ and clothes ID $c_{i}$, we uniformly sample $K$ different clothes IDs, $c_{j}\sim\mathcal{U}(C),\forall j\in\{1,...,K\}$ where $j\neq i$. For a sampled clothes ID $c_{j}$, we obtain its holistic clothes description $R^{c_{j}}$ by leveraging LLaVA and LLaMA (Sec. 2). For a given image $x_{i}$, we get the ID-preserved masked image $\bar{x}_{i}$ by applying its corresponding human-parsed semantic mask, $m_{i}$. To generate $K$ diverse images of the same subject (person $s_{i}$) corresponding to the sampled $K$ clothes IDs, we utilize the extracted clothes descriptions $\{R^{c_{j}}\}_{j=1}^{K}$ and masked image $\bar{x}_{i}$ as inputs to a pretrained Stable Diffusion model [42] for ID-preserving inpainting. The masked image $\bar{x}_{i}$ is used in the forward diffusion process (Eq. (1)), to preserve ID-related information. The reverse denoising process (Eq. (3)) is conditioned on the prompt $R^{c_{j}}$ to generate the clothes-changed images $\{\hat{x}_{ij}\}_{j=1}^{K}$ of subject $s_{i}$ in $x_{i}$, as follows:
@@
We do not train/fine-tune the pretrained diffusion model, but solely use it in an inference setting, thus keeping computational requirements low with a throughput of roughly one generated image per second. Thus, we construct a set of generated data $G_{train}=\{(\hat{x}_{ij},s_{i},c_{ij})_{j=1}^{K}\}_{i=1}^{M}$ for each aforementioned CC-ReID dataset which can be directly utilized to enrich the original training data set, $D_{train}$. It is this data, containing over $2.1$M generated images in total across multiple CC-ReID datasets, that we release for public use as summarized in Table 1.

### v2

@@ Abstract
With the recent exhibited strength of generative diffusion models, an open research question is if images generated by these models can be used to learn better visual representations. While this generative data expansion may suffice for easier visual tasks, we explore its efficacy on a more difficult discriminative task: clothes-changing person re-identification (CC-ReID). CC-ReID aims to match people appearing in non-overlapping cameras, even when they change their clothes across cameras. Not only that current CC-ReID models are constrained by the limited diversity of clothing in current CC-ReID datasets, but generating additional data that retains important personal features for accurate identification is a current challenge. To address this issue we propose DLCR, a novel data expansion framework that leverages pretrained diffusion and large language models (LLMs) to accurately generate diverse images of individuals in varied attire. We generate additional data for five benchmark CC-ReID datasets (PRCC, CCVID, LaST, VC-Clothes, and LTCC) and increase their clothing diversity by $10$$\times$, totaling over $2.1$M generated images. DLCR employs diffusion-based text-guided inpainting, conditioned on clothing prompts constructed using LLMs, to generate synthetic data that only modifies a subject’s clothes, while preserving their personally identifiable features. With this massive increase in data, we introduce two novel strategies – progressive learning and test-time prediction refinement – that reduce training time and boost CC-ReID performance. We validate our method through extensive ablations and experiments, showing massive improvements when training previous CC-ReID methods on our generated data. On the PRCC dataset, we obtain a large top-1 accuracy improvement of $11.3\%$ by training CAL, a state-of-the-art (SOTA) method, with DLCR-generated data. We publicly release our code and generated data for each dataset here: https://github.com/CroitoruAlin/dlcr.
@@ 1 Introduction
In computer vision research, improved or novel model architectures for discriminative tasks are commonly proposed and shown to learn better visual representations than their predecessors. Discriminative methods operate under the assumption that their training datasets are sufficiently large and diverse, with the limiting factor being the model’s ability to learn the desired objective. However, when there is a lack in diversity in the commonly used datasets for some task, the limiting factor in this case becomes the training data itself. Collecting additional in-domain data can be both difficult and expensive, and generative models have only recently become viable enough to generate additional, high-fidelity training data (which we refer to as generative data expansion). Generative data expansion has been previously explored for supervised or self-supervised learning on easier discriminative tasks, such as ImageNet classification [3, 50]. While there are other unexplored, more difficult discriminative visual tasks that may benefit even more from generative data expansion, the difficulty lies in generating accurate, in-domain data. In this work, we tackle this problem in the context of CC-ReID, which is difficult in nature and, due to its relative infancy, a majority of datasets suffer from only containing a small number ($2$-$5$) of clothing outfits per subject [52, 58, 39, 15, 45].
@@
In Person Re-identification [60, 61], the primary objective is to match an individual across images or videos captured by cameras at different times and locations. This task becomes notably challenging due to disparities in the body poses, lighting, backgrounds, viewpoints, and potential occlusions present across cameras. Traditional Re-ID approaches [5, 8, 38], also known as short-term Re-ID, assume that individuals are observed for relatively brief durations and that their attire remains constant throughout videos. However, these assumptions often fall short of practical real-world scenarios. For instance, people’s clothes can change when captured across different camera feeds at different times and days. Consequently, a new branch of Re-ID, known as clothes-changing Re-ID [39, 15, 24, 17], has recently emerged to reliably identify people even when their clothing constantly changes over extended periods of time. Recent CC-ReID works commonly make discriminative modifications to traditional Re-ID model architectures and loss formulations to disentangle structural (clothes-independent) and appearance information (clothes-dependent) in the learned feature space. However, these discriminative approaches are inherently constrained by the aforementioned low clothing variations present in current CC-ReID datasets.
@@
Figure 1: Best viewed with zoom. (Left) Comparing alternative data expansion approaches with our method. Simply changing the color of the clothes using our clothing masks preserves the subject’s ID information, but lacks diversity. Using standard diffusion without inpainting can sometimes introduce diversity, but the ID-related information in an image is destroyed. Our method strongly increases clothing diversity, even with specific and difficult prompts, and preserves the subject’s ID-related information for effective CC-ReID. (Right) DLCR significantly increases the size of training data for five benchmark CC-ReID datasets, generating over $2.1$M diverse, clothes-changed images.
@@
To address these issues, in this work we propose an effective data expansion framework for CC-ReID (named DLCR) that increases the number of clothing variations in CC-ReID datasets by intelligently harnessing generative AI and foundation models (Diffusion and large Language models for CC-ReID). DLCR consists of two stages: (1) Data Generation and (2) Re-ID Training and Prediction Refinement. In the first stage, our objective is to enrich a given CC-ReID training dataset by generating additional samples where clothing items are artificially changed through inpainting. With more diverse training images of a particular person in different clothes, a CC-ReID model will learn better clothes-invariant person features during training. To achieve this objective, we utilize diffusion models [20, 42, 6, 36], which are capable of generating realistic images and inpainting for image editing. However, diffusion models often struggle with preserving the intricate details of a person’s identity during image generation (see Fig. 1). We overcome this by leveraging a human parsing method [26] to produce binary masks that mark only the clothing regions in an image. We then use this binary mask to retain ID-specific portions in an image during diffusion inpainting, such as the face, hair, and body shape, thereby only augmenting the subject’s clothes. Furthermore, existing CC-ReID datasets only use scalar values for each unique clothes ID and lack textual descriptions of specific clothing items, preventing the use of text-guided diffusion models for data generation. Thus, to overcome this issue, we generate explicit clothing descriptions by jointly leveraging a large visual-language model (LLaVA [32]) and a large language model (LLaMA [51]).
@@
In the second stage of DLCR, we propose two effective strategies in which our generated data can be sophisticatedly utilized during CC-ReID training and testing. Firstly, we train various CC-ReID models using our generated data from stage 1 and show how our enriched training data improves performance in clothes-changing settings. To accommodate both the large increase in data and the higher clothing variety of the generated samples, we employ a progressive learning strategy during training, where we gradually introduce the generated clothing variations of a person. This allows for training on 10×10$\times$10 × more data without drastically increasing training time, as well as further assists the model in learning better clothes-invariant person features. During inference, we apply our prediction refinement strategy which is specifically tailored for a CC-ReID retrieval task. We create different variations of a given query by artificially changing the subject’s attire, then ensemble their retrieved similarity scores from the gallery to obtain a refined final prediction.
@@ 2 Related Works
In standard Re-ID, clothing is considered a static attribute, which leads standard Re-ID models to erroneously couple clothing items with specific subjects. CC-ReID is a more difficult paradigm that requires models to intrinsically learn person-specific features that are robust to covariate information, especially clothing items. To achieve this robustness, both discriminative [19, 16, 21, 58, 28, 11, 62, 31, 56, 18] and generative [35, 55, 13, 34, 63, 57] approaches for CC-ReID have been explored. Discriminative approaches focus on elaborate model architectures [30, 15] and loss formulations [19, 16, 21], while generative methods aim to synthesize data that increases the quality or diversity of an existing training set. The common goal of these methods is to enforce the learning of clothes-invariant person features, thereby improving CC-ReID performance on datasets in which subjects are captured in multiple different clothing outfits.
@@
Generative methods: With the aforementioned lack of such diverse datasets, the previously dominant approach has been to use Generative Adversarial Networks (GANs) [14] to synthesize additional training data for specific learning objectives, such as pose-invariance [35, 55, 13], clothes-invariance [34, 63, 57], and improving general performance [64, 29, 23]. However, these GAN-based works also suffer from constraints such as limited diversity/mode collapse [63, 33] and the limited capability of only transferring existing styles from one subject to another [9, 65, 66]. The most important aspect is that GANs are dataset-specific, meaning they require separate training on each dataset before data generation. Furthermore, one of the most successful GAN-based works [63] does not directly use its generated data during training, but rather transfers its dataset-specific generator encoder to the discriminator to produce better clothes-invariant person features, reducing generalizability. On the other hand, diffusion models [36] have recently overtaken GANs in a multitude of other vision domains [10, 6, 43, 44] as a superior class of generative models, yet have not been applied to the CC-ReID domain. Furthermore, the previously discussed limitations of GANs can be overcome by large-scale, pretrained diffusion models which have shown impressive zero-shot generation abilities on a multitude of out-of-domain tasks [40, 42, 37]. DLCR is the first generative approach to harness the benefits of diffusion models for CC-ReID, and consequently, uses the same generator for every dataset. In a novel fashion, we leverage a text-conditioned diffusion inpainting model to synthesize completely new clothing outfits onto any subject in a dataset, separating the process between data generation and discriminative CC-ReID training.
@@ 3 Proposed Approach
DLCR is a generic data expansion method that can be applied to any CC-ReID dataset, wherein the given training data is enriched to improve data diversity for better learning of clothes-invariant features. Our method contains two stages: Data Generation (Sec. 2) and Re-ID Training and Prediction Refinement (Sec. 3.2). In stage 1, DLCR utilizes semantic human masks, a visual language model (LLaVA), a large language model (LLaMA), and text-conditioned diffusion inpainting to controllably change a subject’s clothes in an image. In stage 2, we leverage our generated data, in conjunction with our progressive learning strategy, to train CC-ReID models with 10×10$\times$10 × more data with negligible impacts to training time. We also propose a test-time prediction refinement strategy specifically tailored for CC-ReID retrieval. Before explaining the method in detail, we provide a brief background on diffusion models.
@@ 3.1 Stage 1: Data Generation
In CC-ReID, the training dataset $D_{train}=\{(x_{i},s_{i},c_{i})\}_{i=1}^{M}$ is a collection of $M$ samples, where the $i^{th}$ sample is a triplet consisting of an RGB image $x_{i}$, subject (person) ID $s_{i}$, and clothes ID $c_{i}$. Generating images of a particular subject, $s_{i}$, in different clothing outfits can enrich the training data by improving its diversity. Specifically, in this stage, we generate multiple variants of a training image, $x_{i}$, by artificially inpainting the clothes of subject $s_{i}$. We achieve this objective by using a text-conditioned diffusion inpainting model (Stable Diffusion Inpainting [42]), where text prompts describe the outfits to be generated in the clothing regions specified by a human-parsed semantic mask. These components (also shown in Fig. 2) are described below.
@@
Semantic Mask & Text Extraction: While generating different images of a subject, it is important to preserve the person-relevant information. However, diffusion models struggle to retain this information due to the varied and complex nature of human faces and body shapes. Thus, we apply a semantic human parsing method [26] to each image, $x_{i}$, to obtain an ID-preserving binary mask, $m_{i}$, where only the upper-clothes, lower-clothes, and footwear regions are marked for inpainting. This process is computationally negligible, with throughput as high as 60 masks per second. This allows the person ($s_{i}$) features, such as the face, hair, and general body structure, to be preserved in the generated samples, even after clothes inpainting—an especially critical aspect for enriching CC-ReID data. In general, the set of images belonging to a clothes ID, $c$, can be represented by $X^{c}=\{x^{c}_{i}\}_{i=1}^{n}$. After applying ID-preserving semantic masks to these images, we obtain a set of masked images $\bar{X}^{c}=\{x^{c}_{i}\circ m_{i}\}_{i=1}^{n}$. While CC-ReID datasets provide clothes IDs, they are usually simple scalar values that do not contain descriptive information as to what specific clothing items are present in an image. To extract these additional clothing descriptions for each clothes ID in $D_{train}$, we utilize LLaVA [32] and LLaMA [51] (see Fig. 2). A naive approach would be to simply generate descriptions of clothing items using an LLM, e.g. LLaMA, or to create random clothing descriptions. While this is likely to increase the diversity of the dataset, and consequently, the generalization capacity of the downstream CC-ReID models, it does not alleviate dataset-specific biases. We use LLaVA in order to obtain descriptions of clothing items that are present in the dataset, aiming to reduce the dataset-specific bias (see supplementary). This forces the CC-ReID models to focus on identity features, and ignore clothing features.
@@
Diffusion-Based Clothes Inpainting: A text-conditioned diffusion inpainting model, paired with our extracted clothing descriptions and semantic masks, can be used to controllably generate diverse, synthetic samples. To this end, we propose, for the first time, the use of a diffusion inpainting model for CC-ReID. We first construct a set $C=\{c_{i}\}_{i=1}^{L}$ containing all the unique clothes IDs in the training data $D_{train}$. For each training image $x_{i}$ with subject ID $s_{i}$ and clothes ID $c_{i}$, we uniformly sample $K$ different clothes IDs, $c_{j}\sim\mathcal{U}(C),\forall j\in\{1,...,K\}$ where $j\neq i$. For a sampled clothes ID $c_{j}$, we obtain its holistic clothes description $R^{c_{j}}$ by leveraging LLaVA and LLaMA (Sec. 2). For a given image $x_{i}$, we get the ID-preserved masked image $\bar{x}_{i}$ by applying its corresponding human-parsed semantic mask, $m_{i}$. To generate $K$ diverse images of the same subject (person $s_{i}$) corresponding to the $K$ sampled clothes IDs, we utilize the extracted clothes descriptions $\{R^{c_{j}}\}_{j=1}^{K}$ and masked image $\bar{x}_{i}$ as inputs to a pretrained Stable Diffusion model [42] for ID-preserving inpainting. The masked image $\bar{x}_{i}$ is used in the forward diffusion process (Eq. (1)), to preserve ID-related information. The reverse denoising process (Eq. (3)) is conditioned on the prompt $R^{c_{j}}$ to generate the clothes-changed images $\{\hat{x}_{ij}\}_{j=1}^{K}$ of subject $s_{i}$ in $x_{i}$, as follows:
@@
We do not train/fine-tune the pretrained diffusion model, but solely use it in an inference setting, thus keeping computational requirements low with a throughput of roughly one generated image per second. Thus, we construct a set of generated data $G_{train}=\{(\hat{x}_{ij},s_{i},c_{ij})_{j=1}^{K}\}_{i=1}^{M}$ for each aforementioned CC-ReID dataset, which can be directly utilized to enrich the original training data set, $D_{train}$. It is this data, containing over $2.1$M generated images in total across multiple CC-ReID datasets, that we release for public use, as summarized in Table 1.